When x = 3 and y = 5, by how much does the value of 3x2 – 2y exceed the value of 2x2 – 3y ? i cant remeber how to do these
Answer by stanbon(75887)

When x = 3 and y = 5, by how much does the value of 3x^2 – 2y exceed the value of 2x^2 – 3y ? i cant remeber how to do these
---------------------------------------------
3x^2-2y = 3(3^2)-3(5) = 27-15 = 12
---
2x^2 - 3y = 2(3)^2-3(5) = 18-15 = 3
-------
Ans: 12-3 = 9
==========================
Cheers,
Stan H.